OpenStudy (anonymous):

Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4.

OpenStudy (anonymous):

this is my answer f(x) x 8 1 16 2 24 3 32 4

OpenStudy (anonymous):

Are you asking whether its correct?

OpenStudy (anonymous):

yes lol is my work correct?

OpenStudy (anonymous):

If not can you explain :/

OpenStudy (anonymous):

yep its right.

OpenStudy (anonymous):

it also say to Make a table with integer values of x from 1 to 4.

OpenStudy (anonymous):

so how would the table look? :O

OpenStudy (anonymous):

y(1)=8*1 y(2)=8*2=16 y(3)=8*3=24 y(4)=8*4=32 f(x) x 8 1 16 2 24 3 32 4 It should look like this?

OpenStudy (anonymous):

I'm right anyone ?help what is the answer

OpenStudy (ash2326):

@Esmy After first hour 8 systems got affected after second hour each of the 8 will affect 8 more so=>\(8\times 8\) so it's actually \(y=8^x\), x represents no. of hours x y 1 \(8^1\) 2 \(8^2\)=> 8*8 3 \(8^3\)=>8*8*8 and so on

OpenStudy (anonymous):

Ohhh... So then thats the table I know what I did wrong.

OpenStudy (ash2326):

you got it now?

OpenStudy (anonymous):

I think it would be something like exponential growth, I agree with @ash2326

OpenStudy (anonymous):

yes i got it :)

OpenStudy (anonymous):

thanks for the explanation to help me understand

OpenStudy (ash2326):

You're welcome @Esmy Yeah it's exponential growth @imakebots